### v3

# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

Abstract
Per- and polyfluoroalkyl substances (PFAS), chemicals found in products like non-stick cookware, are unfortunately persistent environmental pollutants with severe health risks. Accurately mapping PFAS contamination is crucial for guiding targeted remediation efforts and protecting public and environmental health, yet detection across large regions remains challenging due to the cost of testing and the difficulty of simulating their spread. In this work, we introduce FOCUS, a geospatial deep learning framework with a label noise-aware loss function, to predict PFAS contamination in surface water over large regions. By integrating hydrological flow data, land cover information, and proximity to known PFAS sources, our approach leverages both spatial and environmental context to improve prediction accuracy. We evaluate the performance of our approach through extensive ablation studies, robustness analysis, real-world validation, and comparative analyses against baselines like sparse segmentation, as well as existing scientific methods, including Kriging and pollutant transport simulations. Results and expert feedback highlight our framework’s potential for scalable PFAS monitoring.
Introduction
PFAS or per- and polyfluoroalkyl substances are persistent “forever chemicals” widely used in industrial and consumer products such as non-stick cookware, waterproof textiles, and firefighting foams (National Institute of Environmental Health Sciences n.d.). Unfortunately, these substances resist degradation and accumulate in water (Langenbach and Wilson 2021; Schroeder, Bond, and Foley 2021), soil (Crone et al. 2019), and living organisms (Environmental Working Group 2023b), with 97% of Americans exhibiting detectable levels in their blood (CDC 2024). Such pervasive exposure is linked to severe health risks including cancers, liver damage, and developmental disorders (Manz 2024). With the high cost of measuring PFAS concentrations in the environment, which leads to sparse ground truth data (Fig. 1), we are left uncertain about targeted remediation to best protect health.
Figure 1: Illustration of limited PFAS contamination data (U.S. EPA, OW 2015)
Figure 2: Left: Classical ML techniques aggregate surrounding pixel data to predict contamination at specific points (depicted as x, y coordinates); Right: DL methods process raster images directly to generate dense PFAS contamination maps in a single pass.
Artificial intelligence (AI) presents a promising avenue for addressing complex health and environmental issues, including identifying PFAS hotspots. In various domains, AI has proven effective at automating labor-intensive tasks and synthesizing domain expertise, whether in land cover mapping (Robinson et al. 2020), agricultural optimization (Kerner et al. 2024), or disease prediction (Bondi-Kelly et al. 2023). In the domain of PFAS contamination prediction, scientists have predominantly applied machine learning (ML) models such as random forests (Breiman 2001) and XGBoost (Chen and Guestrin 2016). For example, (DeLuca et al. 2023) uses a random forest model to predict PFAS contamination via a tabular approach that aggregates environmental features within a 5 km buffer around sample points (Fig. 2). Building on these efforts, we frame PFAS prediction as a geospatial deep learning (DL) task. Our approach can take satellite image formats (i.e., rasters) as input directly, preserving spatial dependencies, reducing the need for extensive feature engineering and aggregation, and leading to more efficiency for ongoing monitoring efforts.
Our contributions in this work are summarized as follows:
  • We propose FOCUS (Framework for EnvirOnmental Contamination with Uncertainty Scaling), a framework with a geospatial DL model for segmentation of PFAS contamination using image formats (rasters) directly.
  • We account for the limited labeled data by developing a novel noise-aware loss, leveraging domain expertise on PFAS spread to fill data gaps with confidence weighting.
  • We benchmark our sparse segmentation framework against other AI models (Hua et al. 2022), tabular ML approaches (DeLuca et al. 2023), and popular methods like pollutant transport simulations (SWAT n.d.) and Kriging (GISGeography 2017), demonstrating our framework’s scalability and accuracy for large-scale PFAS prediction.
  • Through ablation studies, we assess our FOCUS loss function and hyperparameter selection.
We develop this framework together with an environmental nonprofit organization, academic researchers specializing in water quality models and environmental chemistry, and an advisory group with representatives from other nonprofits and agencies, ensuring that our approach is grounded in the latest expertise and real-world environmental concerns.
Related Work
We now summarize existing research about PFAS, spanning analytical methods from scientific domains, to emerging computational approaches, including AI techniques. We also discuss AI techniques for geospatial and sparse data.
Current Ways to Measure and Predict PFAS.
The measurement of PFAS relies on laboratory-based analysis of water, fish tissue, and human blood samples with techniques such as liquid chromatography-tandem mass spectrometry (LC-MS) (Shoemaker and Tettenhorst 2020). PFAS monitoring efforts have made use of such measurement techniques and beyond. For example, states such as Colorado have implemented water testing, treatment grants, and regulatory measures since 2016 (Colorado Department of Public Health and Environment n.d.). In addition, there are community-based water sampling initiatives such as volunteer sampling by the Sierra Club (Sierra Club n.d.) and focused efforts by organizations like the Ecology Center (EcoCenter n.d.). A recent study further underscores the importance of these efforts, revealing associations between PFAS contamination in drinking water and higher COVID-19 mortality rates in the U.S. (Liddie et al. 2024). Resources like the PFAS-Tox Database (Pelch et al. 2022) also assist communities in monitoring PFAS. While highly accurate and important, laboratory-based PFAS measurements are costly, time-intensive, and difficult to scale spatially (Doudrick 2024).
Various modeling approaches have been explored to predict contaminant distribution. Hydrological models such as SWAT (SWAT n.d.) and MODFLOW (Torcoletti 2012) simulate pollutant transport using environmental parameters like flow rates, land use, and weather conditions. While effective, these models require extensive data and are computationally expensive at large scales (Zhi et al. 2024). Kriging, a geostatistical interpolation technique, has been widely used for mapping pollutants, including in soil (Largueche 2006) and groundwater (Singh and Verma 2019). Though useful in data-sparse environments, Kriging relies on simplifying assumptions such as spatial continuity and stationarity (GISGeography 2017) that may not hold for complex contaminants like PFAS, whose transport dynamics and uncertainties are more intricate (see section Results). Complementing these efforts, Salvatore et al. (Salvatore et al. 2022) proposed a “presumptive contamination” model that estimates PFAS presence based solely on proximity to known PFAS-using or -producing facilities. We aim to contribute to monitoring and modeling efforts by providing scalable, data-driven PFAS predictions.
AI Approaches for PFAS Prediction.
Recent advances in AI and geospatial modeling have opened new avenues for tackling the scalability challenges of PFAS detection. For example, DeLuca et al. (DeLuca et al. 2023) used random forests to predict PFAS contamination in fish tissue in the Columbia River Basin, leveraging geospatial data from industrial and military facilities. On a national scale, the USGS applied XGBoost to forecast PFAS occurrence in groundwater (Tokranov et al. 2023) using PFAS sources, groundwater recharge, etc. Such approaches require extensive manual feature engineering, such as aggregating data spatially before prediction is possible (Fig. 2). In addition to being resource-intensive (see section Results), this preprocessing loses spatial context.
Geospatial Models.
Recent advances in geospatial deep learning have introduced foundation models trained on large-scale satellite datasets for improved spatial prediction. Models like Prithvi (Blumenfeld 2023) and SatMAE (Cong et al. 2023) use masked autoencoders (MAE) to learn generalizable spatial representations in a semi-supervised manner, leveraging both labeled and unlabeled data. Unlike traditional ML approaches that require extensive feature engineering (e.g., Fig. 2), these models process raster (image) data directly, preserving spatial dependencies. However, many of these approaches assume that key information is directly observable in satellite imagery, which is not the case for PFAS pollution (see section Results). We therefore adapt Prithvi to use preprocessed, proxy features, such as land cover.
Addressing Data Scarcity.
In environmental data analysis, limited ground truth often necessitates pseudo- or augmented labels, which can introduce noise and lead to overfitting (Safonova et al. 2023). Various strategies have been proposed to mitigate these issues, including label smoothing and robust loss functions (e.g., Huber loss (Gokcesu and Gokcesu 2021), Generalized Cross-Entropy (Zhang and Sabuncu 2018), bootstrapping loss (Reed et al. 2015)), and focal loss (Lin et al. 2018)). Among these, focal loss is particularly effective for addressing class imbalance by down-weighting well-classified examples and focusing the model’s learning on hard, uncertain samples. Weakly supervised methods like the FESTA loss (Hua et al. 2022) further aim to capture spatial and feature relationships. FESTA is specifically designed for segmentation of remote sensing images using sparse annotations assuming spatial continuity, which may not hold for PFAS (see section Results). Furthermore, our incomplete understanding of PFAS-specific hydrological processes hinders the application of both transfer and active learning approaches, making it difficult to estimate uncertainty and find reasonable proxies (Guelfo et al. 2021; Russo et al. 2020). These challenges underscore the need for methods that integrate domain-specific knowledge directly into the training process.
Methodology
This study aims to predict PFAS presence to inform environmental decision-making. In particular, we perform segmentation on surface water area, classifying them as having PFAS concentrations above or below established health advisory thresholds using geospatial data. We next describe FOCUS, including the data preparation pipeline, and the modeling and evaluation approaches.
Figure 3: Overview of dataset curation pipeline
Dataset
To contextualize our data preparation pipeline, Fig. 3 illustrates how ground truth samples are integrated with geospatial data to create the final dataset.
Ground Truth Samples
We use a combination of datasets and health advisories from government agencies in the United States, for both water and fish tissue samples. These are summarized in Table 1 (details in supplement).
Type Use Data Source Total Size
Fish Train/Test EPA (U.S. EPA, OW 2015; US EPA 2024) 866
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy 2025) 114
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy 2023) 293
Water Validation Sampled by Authors 8
Table 1: Datasets used for training, testing, & real-world validation.
We utilize both fish and water data, as they provide snapshots of PFAS contamination with different contexts. For example, fish accumulate PFAS over time, serving as longer-term indicators of water contamination and providing insights into potential exposure risks to anglers (Environmental Working Group 2023a), whereas water sampling provides a measure of PFAS contamination at a specific place and time. Our analysis primarily uses fish data due to the nationally representative extent.
Furthermore, PFAS concentrations are measured as continuous values in these samples, with individual measurements for different species of PFAS. We binarize PFAS concentration measurements into above (1) or below (0) health advisory thresholds, such as the EPA’s Fish and Shellfish Advisory Program health thresholds (U.S. EPA 2024). In particular, we use the cumulative metric called the Hazard Index, which assesses combined risk from individual PFAS compounds by summing their concentration-to-threshold ratios (U.S. Environmental Protection Agency 2023). Note that for fish train/test data, this yields 775 above-threshold (89.5%) and 91 below-threshold (10.5%) samples, reflecting a critical real-world imbalance. For completeness, we also experimented with an alternative labeling scheme using three classes (see supplement).
Raster Image Generation
To spatially integrate the data for training and testing, multi-channel raster images with were generated around each sample point. Each image is a $P\times P$ pixels patch with 30-meter resolution, centered on the geographic coordinates of a sample point, where the optimal value of $P$ is determined in Section 3.3. Once trained, the model can make predictions at any location simply by extracting and processing the corresponding raster, which is done for real-world validation.
These raster images integrate several key features. First, we incorporate readily available rasters such as the National Land Cover Database (NLCD) (U.S. Geological Survey 2024), and flow direction rasters that capture hydrological connectivity using the D8 algorithm in ArcGIS Pro (Esri 2024). In addition, we include discharger location data obtained from the U.S. EPA Enforcement and Compliance History Online (ECHO) (U.S. Environmental Protection Agency n.d.) and convert them into distance rasters using a distance transform (SciPy Developers 2019) (Fig. 4).
Figure 4: Example raster channels: (left) land cover, (center) distances from chemical manufacturing industries, and (right) flow direction.
Ground Truth Masks
Our goal is to perform segmentation to predict PFAS contamination in surface water, but we have sparse point-level contamination measurements rather than dense labels across rasters. To generate dense training labels, every surface water pixel in a patch is labeled according to the point-level sample measurement in that patch (i.e., if the sample point indicates PFAS above the safety thresholds, all surface water pixels are labeled 1; if below, they are labeled 0). Non-surface water pixels are assigned a value of 2. This approach extends point-level data across the patch, assuming that nearby areas have similar contamination.
Noise Masks
This assumption may generate inaccurate labels. To quantify the confidence that a pixel’s label (either 0 or 1) is correct after expansion, we generate a domain expert-informed probability of the label correctness for each surface water pixel $i$:
$p_{\text{final}}=$ $\,\alpha_{1}\cdot p_{\text{dischargers}}+\alpha_{2}\cdot p_{\text{landcover}}+\alpha_{3}\cdot p_{\text{sample\_dist}}$
$\,+\alpha_{4}\cdot p_{\text{downstream}}$ (1)
where $p_{\text{dischargers}}$, $p_{\text{landcover}}$, $p_{\text{sample\_dist}}$, and $p_{\text{downstream}}$ represent probabilities derived from (i) proximity to PFAS dischargers, (ii) land cover type, (iii) distance to other sample points, and (iv) downstream flow, respectively. In detail:
  • Proximity to PFAS Dischargers ($p_{\text{dischargers}}$) Proximity to known PFAS dischargers is a strong indicator of contamination risk, as shown in scientific analyses (Salvatore et al. 2022). If a pixel’s label is 1, we apply an exponential decay function of distance, assigning higher probabilities to pixels closer to known dischargers. For pixels labeled 0, we invert this logic so that being near a discharger lowers confidence in the 0 label.
  • Land Cover Type ($p_{\text{landcover}}$) For pixels labeled 1, surrounding urban or built-up areas received higher probabilities, consistent with research linking PFAS to industrial zones (Dimitrakopoulou et al. 2024). Conversely, for 0-labeled pixels, surrounding undeveloped or forested land cover increased confidence in low contamination.
  • Proximity to Other Sample Points ($p_{\text{sample\_dist}}$) Pixels closer to ground truth sample points were assigned higher probabilities using an exponential decay function based on distance to each sample point in the patch.
  • Downstream Flow ($p_{\text{downstream}}$) PFAS may also travel downstream in aquatic systems, making hydrologic flow another critical factor in contamination spread (Pétré et al. 2022). If a pixel lies downstream of a sample point labeled 1, its probability for label 1 increases, and similarly for label 0. This flow direction channel was generated using ArcGIS Pro (ESRI n.d.) based on digital elevation models (DEMs) (Earth Science Data Systems 2024).
Based on discussions with domain experts and experimental validation (see supplement), we set the weights as $\alpha_{1}=0.4$, $\alpha_{2}=0.2$, $\alpha_{3}=0.1$, and $\alpha_{4}=0.3$. Pixels at ground truth sample points labeled 1 are assigned a probability of 1, reflecting complete confidence. However, for sample points labeled 0, we account for additional uncertainty due to certain PFAS compounds having method detection limits (MDLs) that exceed advisory thresholds [^1] In such cases, a compound may go undetected despite being present at a high-risk level. To address this, the corresponding noise mask value is down-weighted in proportion to the ratio of the advisory threshold to the MDL.
Model
Figure 5: Model overview: high-level representation of FOCUS
Our framework, FOCUS, is inspired by the Prithvi architecture (Blumenfeld 2023) and employs a masked autoencoder (MAE) approach for pretraining (Fig. 5). Instead of relying on Prithvi’s pretrained weights, we pretrain our model on derived geospatial data products (e.g., land cover and distance rasters), which we hypothesize offer more contextual information for capturing the environmental nuances critical to PFAS contamination prediction (see section Results). This pretraining uses masked autoencoding with a mean squared error loss. The pretrained encoder is then fine-tuned with a decoder head dedicated to PFAS prediction.
Loss Function
The loss function used in FOCUS combines the focal loss with noise mask weighting as follows:
$\mathcal{L}_{\text{CE}}(p_{i},y_{i})$ $=-\left[y_{i}\cdot\log(p_{i})+(1-y_{i})\cdot\log(1-p_{i})\right],$ (2)
$\mathcal{L}_{\text{focal}}(p_{i},y_{i})$ $=(1-p_{i})^{\gamma}\cdot\mathcal{L}_{\text{CE}}(p_{i},y_{i}),$ (3)
$\mathcal{L}_{\text{{FOCUS}}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\text{focal}}(p_{i},y_{i})\cdot M_{i}.$ (4)
where $N$ is the total number of pixels, $y_{i}\in\{0,1\}$ is the ground truth label for pixel $i$, $p_{i}$ is the predicted probability of the positive class, and $M_{i}\in[0,1]$ is the corresponding noise mask value indicating confidence in the label.
The parameter $\gamma=2$ is used in the focal loss to focus on hard-to-classify examples, as this value showed best performance in previous studies introducing focal loss (Lin et al. 2018). Notably, class weights to address dataset imbalance are directly incorporated into the binary cross-entropy loss ($\mathcal{L}_{\text{CE}}$), removing the need for an explicit alpha term in focal loss. By combining focal loss and noise mask weighting, the loss function addresses label noise and class imbalance while emphasizing high-confidence pixels during training.
Experiments
Baselines
To contextualize our proposed framework for PFAS contamination prediction, we introduce several baseline approaches commonly employed in the field:
  • Pollutant transport simulation: A scientific, process-based approach that models contaminant movement using hydrological principles (SWAT n.d.).
  • Landsat-based method: Uses the original Prithvi weights to predict PFAS from raw satellite imagery (Landsat 7 multispectral (U.S. Geological Survey n.d.)).
  • FESTA loss approach: An alternative state-of-the-art technique designed for sparse segmentation using sparse point data. (Hua et al. 2022).
  • Kriging: A geostatistical interpolation technique (Esri n.d.) that predicts contamination based on spatial relationships between points.
  • Random forest: Following (DeLuca et al. 2023), this model aggregates environmental features within a 5 km buffer around each sample (e.g., calculating the percentage of various land cover types) to predict contamination at individual points.
  • Rule-based approach: Assigns contamination labels based on predefined environmental heuristics, replicating our noise mask approach to compute a contamination probability (but without ($p_{\text{sample\_dist}}$)), which is then thresholded to yield binary predictions.
Training Configuration
We now describe our experimental setup, with more details in the supplement. We conducted experiments per year, allowing us to evaluate the model’s ability to fill spatial data gaps rather than forecast across time. We split each year’s data into 80% for training and 20% for testing, ensuring that training and testing occur within the same year’s distribution. To ensure geographic diversity and avoid spatial leakage, we further ensured that the test set was geographically disjoint from the training set.
A batch size of 4 and a learning rate of $5\times 10^{-4}$ were adopted to preserve the stability of our pretrained backbone, minimizing the risk of large, destabilizing updates. In addition, we employed AdamW (Loshchilov and Hutter 2019) as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We tuned the learning rate and batch size empirically by monitoring training stability. The final values were selected based on consistent convergence and favorable test-time metrics across years. Furthermore, we used a custom learning rate scheduler with an initial warmup phase followed by polynomial decay, which helped stabilize convergence and ensured that the model gradually adapted to the limited data.
Figure 6: Comparison of performance across different methods and years. Results are averaged over 3 random seeds for FOCUS and applicable baselines (Landsat-based, FESTA loss, rule-based).
Ablation Study
We conduct ablation studies across the following dimensions:
  • Noise Weights: We evaluate the impact of incorporating noise masks into the FOCUS loss function. With noise masks, pixel-wise confidence scores scale focal loss to emphasize high-confidence pixels; without them, the loss reduces to standard focal loss (Lin et al. 2018).
  • Raster Size: To understand the influence of spatial context, we compare raster patches of $512\times 512$ and $256\times 256$ pixels. Larger patches capture broader context, potentially including contamination sources farther from the sample point, while smaller patches reduce label noise and focus on localized features.
While the primary focus of this work is on PFAS contamination, we also explored the generalizability of our framework to land cover classification, and robustness through a consistency analysis detailed in the supplement.
Metrics
The reported metrics include accuracy, Intersection over Union (IoU), F-score, precision, and recall. Performance metrics in this study are reported exclusively from evaluations using only ground truth samples. In addition, we assess timing efficiency as a metric to evaluate the model’s computational scalability.
Results
FOCUS Outperforms Existing Baselines
Main Results
We first compared FOCUS against six alternative approaches for PFAS contamination prediction as shown in Fig. 6. Overall, FOCUS achieves better performance across all the years and most metrics, aside from accuracy in 2019 and 2022, and precision in 2022. The Landsat-based method yielded moderate results, suggesting that geospatial data products are more effective for capturing PFAS contamination patterns than relying solely on raw satellite data. The random forest approach, despite achieving acceptable accuracy, suffered from lower F-score, precision, and recall, likely due to the limited spatial context from per-point feature aggregation. Similarly, although Kriging provided decent accuracy and precision, it struggled with recall, indicating difficulties in detecting all instances of contamination, and pollutant transport simulation demonstrated weak overall performance. The FESTA loss approach showed suboptimal performance in our experiments. This may be attributed to our dataset’s exclusive use of point data and the complex dynamics of PFAS contamination, which may challenge the spatial continuity assumptions of the FESTA loss. While rule-based masks, designed to approximate contamination using domain-specific heuristics only, can provide a reasonable estimate (and prove useful for computing noise masks, see section Methodology), they fall short in accuracy compared to the data-driven approach. This finding underscores the importance of incorporating domain expertise and AI for more reliable PFAS prediction.
Time Efficiency
We followed the methodology of DeLuca et al. (DeLuca et al. 2023), which uses similar environmental features, dataset, and a random forest model, as a baseline for efficiency comparisons. Table 2 compares feature extraction and inference times between the ML method and our FOCUS approach over a 1 km2 area and in Northern Michigan (NM) (approximately 44,000 km2).
For feature extraction, the ML baseline required aggregating environmental features within a 1 km buffer per point, taking an average of 1.2 minutes per point and nearly 2 days over the NM region. In contrast, FOCUS processes full 1 km2 raster patches in 4.7 seconds on average, reducing extraction time to 3.2 hours for NM. Inference times were comparable: 7.5 seconds for ML vs. 13 seconds for FOCUS over NM. Table 2 summarizes these results. All times were averaged over 10 runs. Feature extraction was performed on an Intel Xeon Gold 6154 CPU (3.0 GHz), and inference used an AMD Ryzen Threadripper PRO 7985WX with an NVIDIA RTX 6000 Ada GPU.
Feature extraction Inference
ML DL ML DL
For 1 km2 area 1.2 $\pm$ 0.2 mins 4.7 $\pm$ 0.5 sec 0.0003 $\pm$ 0.0002 sec 0.0005 $\pm$ 0.0001 sec
For NM 2 days 3.2 hrs $\pm$ 0.4 hrs 7.5 $\pm$ 1 sec 13 $\pm$ 2 sec
Table 2: Comparison of performance for feature extraction and inference using random forest and FOCUS over 1 km2 area and Northern Michigan (NM).
FOCUS Applies in the Real World
We next conducted two complementary validation efforts of predicted contamination maps of new regions – beyond regions surrounding ground truth sample points in the train and test sets – using field-collected data (see Table 1). First, our research team collected surface water samples from multiple local sites in Michigan. To reiterate, these represent previously unsampled regions and provide an independent means of assessing model generalization beyond our test set. Notably, as shown in Fig. 7, the model correctly predicted 8 collected points in this evaluation. Technical details on sample analysis are available in the supplement.
Figure 7: Real-world validation using surface water samples collected by our team. Newly sampled locations (green) are overlaid on model predictions, with red indicating areas of high predicted contamination, and dark green indicating areas of lower predicted contamination.
Second, we evaluated our 2022 fish tissue model on the publicly available MPART fish dataset from 2022 (Michigan Department of Environment, Great Lakes, and Energy 2025), which includes PFAS measurements in fish tissue from across Michigan, collected independently from our fish train and test sets. Results are comparable to Fig. 6.
Metric Accuracy IoU F1 Precision Recall
MPART Fish Samples 84 $\pm$ 2 % 61 $\pm$ 1 % 74 $\pm$ 1 % 74 $\pm$ 1 % 73 $\pm$ 2 %
Table 3: FOCUS performance on fish tissue samples collected independently of training data. Results averaged over 3 random seeds.
FOCUS Noise Masks Improve Performance
2008 (%) 2019 (%) 2022 (%)
Metric Focal only FOCUS Focal only FOCUS Focal only FOCUS
Accuracy 79 $\pm$ 2 88 $\pm$ 1 89 $\pm$ 1 89 $\pm$ 1 84 $\pm$ 3 86 $\pm$ 2
IoU 56 $\pm$ 3 65 $\pm$ 2 67 $\pm$ 2 71 $\pm$ 2 65 $\pm$ 3 68 $\pm$ 2
F-score 70 $\pm$ 3 77 $\pm$ 2 79 $\pm$ 1 82 $\pm$ 2 77 $\pm$ 3 80 $\pm$ 2
Precision 68 $\pm$ 2 83 $\pm$ 2 78 $\pm$ 2 78 $\pm$ 2 74 $\pm$ 3 76 $\pm$ 2
Recall 74 $\pm$ 2 73 $\pm$ 2 80 $\pm$ 1 89 $\pm$ 1 91 $\pm$ 1 92 $\pm$ 2
Table 4: Ablation showing impact of noise-aware loss (FOCUS) vs. standard focal loss. Results averaged over 3 random seeds. Bold values denote the best result per year.
Table 4 evaluates the effect of incorporating noise masks in the FOCUS loss computation, compared to focal loss only (i.e., removing $M_{i}$ from Equation 4). We find that the addition of noise masks consistently improves performance, though FOCUS may sometimes favor precision over sensitivity. This is due to emphasizing high-confidence pixels and down-weighting uncertain ones during training.
Note that this ablation is conducted at the $256\times 256$ resolution, which was selected based on the ablation study comparing it with $512\times 512$ patches, detailed in the supplement. The smaller resolution consistently performed better, likely due to reduced label noise propagation across larger contexts. We did not evaluate patch sizes smaller than $256\times 256$, as excessively small patches can compromise accuracy and increase variability (Lechner et al. 2009).
FOCUS has Robust Results
2008 (%) 2019 (%) 2022 (%)
Metric Focal Only FOCUS Focal Only FOCUS Focal Only FOCUS
Accuracy 80 $\pm$ 2 85 $\pm$ 1 92 $\pm$ 1 96 $\pm$ 1 90 $\pm$ 2 91 $\pm$ 1
IoU 64 $\pm$ 2 67 $\pm$ 1 65 $\pm$ 2 73 $\pm$ 1 75 $\pm$ 2 77 $\pm$ 1
F-score 77 $\pm$ 1 79 $\pm$ 1 76 $\pm$ 2 82 $\pm$ 1 85 $\pm$ 1 87 $\pm$ 1
Precision 78 $\pm$ 2 87 $\pm$ 1 70 $\pm$ 3 83 $\pm$ 2 81 $\pm$ 2 83 $\pm$ 1
Recall 76 $\pm$ 2 77 $\pm$ 1 96 $\pm$ 1 91 $\pm$ 2 94 $\pm$ 1 92 $\pm$ 1
Table 5: Cross-state generalization results ($\text{mean}\pm\text{std}$). Bold values denote the best result per year.
To further assess the generalizability of our model to geographically unseen regions, we performed a spatial cross-validation experiment by partitioning data based on U.S. states. Specifically, we selected three distinct sets of five states each to serve as held-out test sets. The remaining states in each case were used exclusively for training. This ensures that no spatial overlap exists between training and testing regions, mimicking the real-world scenario where models are deployed in previously unsampled areas. FOCUS again consistently outperforms focal loss alone across all years, highlighting its ability to generalize spatially to unseen regions and its robustness in distinguishing contaminated areas even under strong distribution shifts.
To assess the spatial robustness of our model, we also performed a consistency analysis across all 49 continental U.S. states. This involved comparing predictions over overlapping regions from multiple randomly sampled patches per state, with varying degrees of overlap (56×56 and 156×156 pixels). Agreement was measured as the proportion of matching predicted labels across overlaps. We found high agreement scores (above 93% on average), indicating that our model is locally consistent in its predictions. Full details are provided in the supplement.
Results on model calibration (i.e., Expected Calibration Error (ECE) (Guo et al. 2017)) are provided in the supplement, showing low ECE values ($\leq 0.1$) and indicating that predicted probabilities are well-aligned with observed outcomes.
Collaboration with Stakeholders
To build this model, and ultimately deploy it in a way that meets real-world needs, we are collaborating with a nonprofit environmental organization. This partnership has helped shape both the technical and practical framing of our work. We meet biweekly to discuss results, align on next steps, and ensure that public-facing outputs, such as PFAS contamination maps, are communicated transparently and responsibly. Recognizing the ethical implications of sharing such information, including for public health, environmental justice, and even property values, we are also currently engaging with an advisory group with representatives from other non-profits and agencies. This process will help tailor outputs to community needs and support responsible use in remediation and policy decisions.
Conclusion and Future Work
PFAS contamination is a global environmental issue due to its persistence in water, soil, and wildlife. While lab-based methods like LC-MS offer high accuracy, they are expensive and slow. Alternative approaches, such as spatial interpolation, remote sensing, ML methods, and hydrological models, provide insights, but struggle with feature engineering and loss of spatial context.
Our geospatial DL framework, FOCUS, directly processes multi-channel raster data and integrates a novel noise-aware loss, improving detection performance and preserving spatial dependencies, making it a promising tool for large-scale PFAS contamination mapping that can help identify high-risk areas.
Future work will explore uncertainty estimation to guide targeted sampling and enhance prediction robustness. We also aim to incorporate temporal modeling using multi-year data to track and forecast evolving PFAS patterns. Finally, we plan to build on public contamination maps, such as those from the Environmental Working Group (Environmental Working Group 2024) and (U.S. Geological Survey 2023), by deploying our AI-based framework responsibly.
Supplement
To help navigate the supplementary material, Table 6 provides an overview linking each supplementary subsection to the corresponding sections in the main paper where they are mentioned or expanded upon.
Supplement Subsection Referenced in Main Paper Section(s)
A. Model Calibration Section 4.4
B. Noise Mask Weights Ablation Section 3.1
C. Kriging Section 3.3
D. Pollutant Transport Simulation Section 3.3
E. Consistency Analysis Section 4.4
F. Evaluating the Impact of Noise Masks on Segmentation of Land Cover Types Section 3.3
G. Dataset Section 3.1, Section 3.3
H. FOCUS Per-Class Results -
I. Ternary Classification of Hazard Levels Section 3.1
J. Statistical Significance of Performance Gains -
K. Distance-to-Industry Rasters. Section 3.1
L. Ablation Results for Image Resolution Section 4.3
M. Hyperparameter Tuning and Selection Section 3.3
N. Technical Details on Sample Analysis and Training Data for Real-World Predictions Section 4.2
Table 6: Mapping between supplementary sections and corresponding references in the main paper.
A Model Calibration
We evaluated model calibration using ECE as introduced in Section 4.4, which quantifies the discrepancy between predicted probabilities and observed outcomes where lower scores indicate better calibration. Table 7 shows the ECE scores for our model trained with FOCUS loss for the years 2008, 2019, and 2022.
Metric 2008 2019 2022
ECE Score 0.1 0.1 0.07
Table 7: ECE Scores for Models Trained with FOCUS Loss Across Different Years.
The very low ECE values, that is, 0.1 for 2008 and 2019, and 0.07 for 2022, demonstrate that our model’s predicted probabilities are closely aligned with the observed outcomes, indicating excellent calibration.
B Noise Mask Weights Ablation
Our ablation study on the 2008 PFAS dataset evaluated a comprehensive set of 24 noise mask weight configurations through a grid search over four input types: dischargers, land cover, flow direction, and distance to sample points as mentioned in Section 3.1. For each configuration, we varied the percentage weights assigned to these four sources of uncertainty and measured segmentation performance across multiple metrics. The best-performing configuration assigns 40% weight to dischargers, 20% to land cover, 10% to sample distance, and 30% to flow direction. This setup achieved the highest scores in accuracy, IoU, F-score, and precision, even though a few alternate combinations yielded marginally higher recall. These results align with domain expert guidance that emphasizes the importance of industrial activity in driving contamination patterns. Full performance results for all configurations are reported in Table 8, demonstrating that our selected weighting scheme most effectively balances the different environmental signals for robust PFAS contamination prediction.
Dischargers (%) Landcover (%) Flow Dir. (%) Sample Dist. (%) IoU (%) F-score (%) Precision (%) Recall (%) Accuracy (%)
40 20 30 10 65 77 83 73 88
40 30 20 10 44 52 67 50 76
40 10 30 20 48 60 68 57 80
40 10 20 30 43 51 52 51 78
40 30 10 20 49 59 63 58 82
40 20 10 30 47 55 67 54 84
30 40 10 20 56 69 93 61 80
30 40 20 10 44 58 58 64 69
30 10 40 20 46 56 59 55 76
30 10 20 40 46 60 59 65 71
30 20 40 40 54 67 67 67 82
30 20 10 40 51 63 63 65 78
30 10 30 30 42 46 42 50 84
10 40 30 20 59 71 73 69 85
10 40 20 40 47 59 58 59 76
10 30 40 20 43 51 52 51 78
10 30 20 40 41 57 62 73 62
10 20 40 10 49 63 63 70 73
10 20 30 40 44 55 55 55 76
20 40 30 40 51 67 68 77 69
20 40 10 30 51 61 68 59 84
20 30 40 10 47 62 63 73 69
20 30 20 30 45 60 64 75 65
20 10 40 30 35 52 64 64 47
Table 8: Performance across all noise mask weight configurations using the 2008 PFAS dataset.
C Kriging
We employed Kriging as a spatial interpolation technique to estimate PFAS, as mentioned in Section 3.3. First, we examined the spatial structure of our data by calculating an empirical semivariogram. Using latitude and longitude coordinates alongside the binary target variable, we computed pairwise distances and differences to produce a semivariogram plot, enabling us to observe how variance changed with increasing distance. This step informed the selection of appropriate variogram models (e.g., linear or spherical) for Kriging. Based on this analysis, we selected a spherical variogram model, which best captured the spatial dependencies observed in our data.
Next, we split the data on a state-by-state basis, ensuring geographic diversity between the training and testing subsets. We explored both Ordinary Kriging (assuming a stationary mean across the study area) and Universal Kriging (introducing a drift term to account for regional linear trends), but the two approaches produced comparable results on our dataset. After training each Kriging model on the geographic regions in the training set, we predicted PFAS presence at test set locations. Since we treated PFAS presence as a binary outcome, the resulting continuous Kriging predictions were thresholded at 0.5. We then computed standard classification metrics (such as precision, recall, and F1-score) to evaluate each model’s ability to distinguish between high contamination versus low contamination sites.
While Kriging provided a practical means of interpolating PFAS presence, its reliance on variogram assumptions and spatial stationarity can limit accuracy in areas with highly heterogeneous conditions.
D Pollutant Transport Simulation
As outlined in Section 3.3, instead of running a simulation like SWAT directly, which is designed for general watershed hydrology modeling but lacks established parameterization for PFAS transport, we implemented a streamlined Python-based workflow more tailored to PFAS-related surface water contamination. This approach allowed us to integrate domain-relevant features without relying on assumptions unsupported by current PFAS science.
Figure 8: Example binary output from the pollutant transport simulation: 0 represents low contamination, and 1 represents high contamination. The simulation approximates the distribution of PFAS contamination based on hydrological and environmental parameters, providing a practical alternative to full-scale SWAT simulations.
First, we assign initial pollutant concentrations in each patch by setting cells flagged as “dischargers” to a high baseline value (e.g., 100), while land cover–based default concentrations provide moderately-low to low initial contamination levels for other cells. We then construct a Hydrologic Response Unit (HRU) parameter table that assigns infiltration and runoff values based on a cell’s land cover, soil type, and slope. Using these parameters, each patch’s land cover and soil rasters define infiltration/runoff ratios for every grid cell. We incorporate two key rasters; flow direction and flow accumulation, both exported from System for Automated Geoscientific Analyses (SAGA); a GIS and geospatial analysis tool focused on geospatial data processing, analysis, and visualization, to specify how water and pollutant mass move downstream. In each iteration, a fraction of the pollutant mass in each cell is transferred to its downstream neighbor according to the cell’s infiltration/runoff factors, guided by the flow direction raster and scaled by the flow accumulation values. Repeating this process causes pollutant mass to concentrate in lower-lying or higher-flow cells, thereby modeling how contamination evolves over time within the patch.
After the simulation converges, we produce a final pollutant concentration raster, thresholded by the median value of concentrations across all patches for the year to generate a binary contamination map, as illustrated in Fig. 8. The binary outputs, with values of 0 and 1 representing low and high contamination respectively, visualizes the simulation’s predictions of contamination distribution. These outputs are compared against test set patches containing actual observed PFAS presence in surface water to evaluate how accurately the simulation captures observed contamination patterns. Although this standalone approach remains computationally non-trivial, it is more tractable for batch processing of multiple patches than running a full SWAT project repeatedly. Consequently, it provides a practical baseline for pollutant transport modeling within our broader framework, allowing us to compare simulated outputs to both real-world data and other modeling approaches.
E Consistency Analysis
To evaluate local spatial consistency as highlighted in Section 4.4, we applied a robustness test across all 49 continental U.S. states. For each state, we extracted two overlapping image patches; one with a 56×56 pixel overlap and another with 156×156, using three different random locations to ensure stability. We then compared the model’s predictions over the overlapping regions by computing pixel-wise agreement. Agreement is the proportion of matching predicted labels across overlapping pixels between two patches:
$\text{Agreement}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(y_{i}^{(1)}=y_{i}^{(2)}),$ (5)
where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are the predicted labels for the $i^{\text{th}}$ pixel in each patch, and $N$ is the total number of overlapping pixels.
Table 9 reports the agreement across different years using our final FOCUS model. The high consistency scores across both overlap sizes demonstrate that the model generates stable predictions for the same spatial region even when seen in different contexts. This supports the reliability of our framework in real-world geospatial settings.
Year 56×56 Overlap 156×156 Overlap
2008 95 $\pm$ 1 % 93 $\pm$ 1 %
2019 96 $\pm$ 2 % 98 $\pm$ 2 %
2022 93 $\pm$ 1 % 94 $\pm$ 1 %
Table 9: Spatial consistency measured as prediction agreement over overlapping patch regions.
F Evaluating the Impact of Noise Masks on Segmentation of Land Cover Types
In this experiment, as introduced in Section 3.3, we focus on distinguishing between two pairs of land cover (LC) types:
  • Evergreen Forests vs Barren Land (Class 0: Barren Land, Class 1: Evergreen Forest)
  • Deciduous Forests vs Water (Class 0: Deciduous Forest, Class 1: Water)
In this experiment, we simulate uncertainty in the labeling process to align with the uncertainty potentially present in our PFAS ground truth label expansion. Specifically, we randomly throw out half of the valid class pixels and replace them with the other class’s label. We then generate noise masks using only the NDVI values, which are used during the training process to handle the uncertainty.
Results
Below are the macro-averaged results showing the performance metrics for the Evergreen Forests vs Barren Land and Deciduous Forests vs Water pairs, both with and without noise masks. These LC image patches are aligned with the same spatial footprint as the PFAS patches to ensure consistency in coverage. The evaluation is conducted on at most 2 pixels per image patch, simulating sparse ground truth availability similar to the PFAS setting. We see that including noise masks is useful in this setting as well.
Evergreen Forests vs Barren Land Performance Metrics
Method IoU F1-Score Precision Recall Accuracy
With Noise Masks 0.38 0.55 0.56 0.56 0.55
Without Noise Masks 0.24 0.33 0.74 0.50 0.48
Table 10: Performance Comparison for Evergreen Forests vs Barren Land: Simulated Uncertainty in PFAS Pseudolabeling
Deciduous Forests vs Water Performance Metrics
Method IoU F1-Score Precision Recall Accuracy
With Noise Masks 0.38 0.55 0.55 0.56 0.55
Without Noise Masks 0.33 0.49 0.51 0.51 0.52
Table 11: Performance Comparison for Deciduous Forests vs Water: Simulated Uncertainty in PFAS Pseudolabeling
G Dataset
As introduced in Section 3.3 and expanded upon here, the dataset was curated to ensure that patches included in the training and testing sets were geographically disjoint, avoiding any overlap.
In cases where two or more patches overlapped, as illustrated in Fig. 9, all such patches were assigned to the same set (either training or testing). This ensured that no part of a test patch had been seen during training, preserving the integrity of the evaluation.
Figure 9: Example of overlapping patch assignment. The figure illustrates two slightly overlapping patches overlaid on the U.S. map, each containing distinct sample points. To maintain disjoint training and testing sets, both patches are assigned to the same set (either training or testing), ensuring no part of a test patch overlaps with any patch seen during training.
Additionally, Table 12 expands on the table in Section 3.1 to include class distribution. There is a significant class imbalance that we observe in the real world: unfortunately, a majority of water bodies have PFAS contamination that exceed health guidelines, as found in prior scientific work on PFAS prediction as well  [^2] See: PFAS pollute 83% of U.S. waterways, E&E News by Politico (2023). Available at: https://www.eenews.net/articles/pfas-pollute-83-of-u-s-waterways/ . To affirm that there is in fact signal to our predictions, we provide class-specific results in Section H. For instance, we find approximately 83% F-score for class 0 for the year 2024, showing we are truly predicting class 0, and better than random. We also add a ternary classification in Section I to provide additional levels of risk.
Type Use Data Source Total Size % Pos / Neg
Fish Train/Test EPA (U.S. EPA, OW 2015; US EPA 2024) 866 89.5% / 10.5%
Fish Validation MPART (Michigan Department of Environment, Great Lakes, and Energy 2025) 114 77.2% / 22.8%
Water Train/Test MPART (Michigan Department of Environment, Great Lakes, and Energy 2023) 293 66.9% / 33.1%
Water Real-World Validation Sampled by Authors 8 100% / 0%
Table 12: Datasets used for training, testing, and real-world validation, including total size and class balance.
H FOCUS Per-Class Results
To better understand model behavior across classes, Table 13 reports per-class performance metrics for FOCUS across each year, including results on the 2024 MPART surface water dataset.
2008 (%) 2019 (%) 2022 (%) 2024 (%)
Metric Class 0 Class 1 Class 0 Class 1 Class 0 Class 1 Class 0 Class 1
Accuracy 88 $\pm$ 1 88 $\pm$ 1 89 $\pm$ 1 89 $\pm$ 1 86 $\pm$ 2 86 $\pm$ 2 86 $\pm$ 2 86 $\pm$ 1
IoU 43 $\pm$ 2 87 $\pm$ 2 53 $\pm$ 2 88 $\pm$ 2 52 $\pm$ 3 83 $\pm$ 2 71 $\pm$ 2 78 $\pm$ 2
F-score 60 $\pm$ 2 93 $\pm$ 2 70 $\pm$ 2 93 $\pm$ 2 68 $\pm$ 3 91 $\pm$ 2 83 $\pm$ 2 88 $\pm$ 2
Precision 75 $\pm$ 3 90 $\pm$ 2 57 $\pm$ 2 98 $\pm$ 2 52 $\pm$ 3 99 $\pm$ 1 99 $\pm$ 1 78 $\pm$ 2
Recall 50 $\pm$ 2 96 $\pm$ 2 89 $\pm$ 1 89 $\pm$ 1 99 $\pm$ 1 83 $\pm$ 2 71 $\pm$ 1 99 $\pm$ 1
Table 13: Per-class performance of our model using FOCUS across four years (MPART surface water results at 2024). Results averaged over 3 random seeds.
I Ternary Classification of Hazard Levels
To enable finer-grained prioritization of contamination severity, we extended our binary classification framework to a ternary setting, as highlighted in Section 3.1. Specifically, each fish tissue sample was assigned to one of three classes based on its hazard index (HI), defined as the ratio of measured PFAS concentration to the EPA’s reference dose:
  • Class 0 (Low): HI << 1 (i.e., below the EPA baseline threshold)
  • Class 1 (Medium): $1\leq$ HI $\leq 1000$
  • Class 2 (High): HI >> 1000
[…]
Metric 2008 (%) 2019 (%) 2022 (%)
Accuracy 55 $\pm$ 1 72 $\pm$ 1 65 $\pm$ 1
IoU 38 $\pm$ 1 48 $\pm$ 2 44 $\pm$ 2
F-score 54 $\pm$ 1 63 $\pm$ 1 61 $\pm$ 1
Precision 66 $\pm$ 2 70 $\pm$ 2 60 $\pm$ 2
Recall 65 $\pm$ 2 61 $\pm$ 1 62 $\pm$ 2
Table 14: Performance of FOCUS under ternary classification (0: low, 1: medium, 2: high). Results averaged over 3 random seeds.
Results demonstrate signal, but are less strong compared to FOCUS binary classification performance, given the added complexity of the ternary task.
J Statistical Significance of Performance Gains
To evaluate the robustness of our improvements over the focal-only baseline, we conducted statistical significance testing using the Wilcoxon signed-rank test, a non-parametric paired test appropriate for comparing performance across random seeds.
For each year (2008, 2019, 2022), we collected the F1-scores from multiple runs (5 seeds per year) and compared the FOCUS results against focal-only using a two-sided Wilcoxon signed-rank test. As shown in Table 15, FOCUS achieved statistically significant improvements in all three years ($p=0.03125$), despite the relatively small number of seeds. Notably, for every random seed, the F1-score achieved with FOCUS was consistently higher than the focal-only counterpart.
These results confirm that the performance gains from our noise-aware loss are not due to chance, but are statistically reliable across different random initializations.
Year Wilcoxon Statistic $p$-value
2008 0.0 0.03125
2019 0.0 0.03125
2022 0.0 0.03125
Table 15: Wilcoxon signed-rank test comparing F1-scores of FOCUS and focal-only across 5 random seeds for each year.
K Distance-to-Industry Rasters
To capture potential industrial sources of PFAS contamination, we constructed distance rasters for each relevant facility type using data from the EPA’s ECHO database as mentioned in Section 3.1. For each year (e.g., 2008, 2019, 2022), we identified all facilities that were active on or before that year, ensuring historical context was preserved. Facility types were grouped by industrial category (e.g., airports, landfills, chemical manufacturers), and a separate Euclidean distance raster was generated for each category, where pixel values represent the distance to the nearest facility of that type.
The total number of distance raster channels therefore varies by year, depending on which facility types were present in the ECHO dataset up to that point. For example, a given year may include distance rasters for 15 industry types, while another may include 41.
L Ablation Results for Image Resolution
Table 16 presents the results of our image resolution ablation study, as highlighted in Section 4.3. We compare the model’s performance when trained on $256\times 256$ versus $512\times 512$ input resolutions, without the use of noise-aware training.
2008 (%) 2019 (%) 2022 (%)
Metric 256 512 256 512 256 512
Accuracy 79 $\pm$ 2 75 $\pm$ 3 89 $\pm$ 1 85 $\pm$ 2 84 $\pm$ 3 78 $\pm$ 3
IoU 56 $\pm$ 3 52 $\pm$ 2 67 $\pm$ 2 66 $\pm$ 2 65 $\pm$ 3 56 $\pm$ 2
F-score 70 $\pm$ 3 67 $\pm$ 3 79 $\pm$ 1 78 $\pm$ 2 77 $\pm$ 3 71 $\pm$ 3
Precision 68 $\pm$ 2 66 $\pm$ 3 78 $\pm$ 2 76 $\pm$ 2 74 $\pm$ 3 70 $\pm$ 3
Recall 74 $\pm$ 2 72 $\pm$ 2 80 $\pm$ 1 83 $\pm$ 2 91 $\pm$ 1 84 $\pm$ 2
Table 16: Comparison of performance between $256\times 256$ and $512\times 512$ image resolutions without noise-aware training. Bold values indicate the better result.
M Hyperparameter Tuning and Selection
We tuned the learning rate and batch size empirically by monitoring training stability as outlined in Section 3.3. Specifically, we experimented with learning rates in $\{1\times 10^{-3},5\times 10^{-4},1\times 10^{-4}\}$ and batch sizes in $\{2,4,8\}$. We selected a batch size of 4 and a learning rate of $5\times 10^{-4}$ as they provided the best stability while preserving the benefits of our pretrained backbone, minimizing the risk of large, destabilizing updates. Notably, these final hyperparameter values are consistent with those adopted in the original Prithvi paper, further supporting their suitability in this context.
N Technical Details on Sample Analysis and Training Data for Real-World Predictions
As introduced in Section 4.2, this section provides additional technical details on the real-world PFAS samples collected and analyzed by our team.
Sample Collection. New 250 mL HDPE (High-Density Polyethylene) bottles were triple rinsed with ultrapure water before sample collection in the field.
Sample Preparation. Samples were extracted using solid phase extraction, guided by EPA Method 1633. Each sample was spiked with 200 µL of isotopically mass-labeled standards (Wellington).
Targeted Sample Analysis. Targeted PFAS analysis was conducted using a Thermo Exploris 240 Orbitrap mass spectrometer (MS) coupled with a Vanquish ultra-high-performance liquid chromatograph (UHPLC-MS). A C-18 column was used, and the MS was operated in negative electrospray ionization (ESI-) mode.
Model Used for Prediction. The model used to generate predictions for these newly collected samples was trained on 2024 surface water samples from the Michigan MPART dataset. Given the temporal proximity and coverage of this dataset, it provided a decent foundation for making informed predictions on samples collected in mid-2025.